Recitation 11 MA113:004–006 31 September 1998 Fall 1998

Reminders: 1. A notebook is on reserve in the math library for this course. This notebook contains tests from last year and solutions to tests, homework quizzes and projects. The math library is located in the basement of the Patterson Office Tower.

2. The next homework is due on Wednesday, 7 October. The assignment is §2.3 #12, 20 §2.4 #6, 14.

Below is a selection of problems related to sections 2.4. These problems will not be collected or graded. However, you should understand how to work each of these problems. You should begin working on these problems in groups in recitation. You will probably want to finish these problems outside of class. If you have questions, please ask your TA or instructor. If you find a problem difficult, consider working similar problems from the text for additional practice.

1. (Algebra review) Expand (a + b)^{2} by multiplying out term by term. Many of you know this as the “foil” method.

2. (Algebra review) Expand (a + b)^{3}. To do this, write (a + b)^{3} = (a + b)^{2}(a + b) and use the known result for (a + b)^{2}. Then multiply out all six terms and collect like terms. You may check your answer by looking in the front cover of the text for this course. Does your answer agree with the result of the simple binomial theorem you were asked to prove in recitation 7.

3. (Algebra review) Solve the equation for x.

y = \frac{1}{x + a^{2}} .

4. Work problems #3–6, 11, 13, 17, 29, 33, 43 in §2.4 of Stewart.

5. Finish problems from sections 2.1–2.3.